OpenStudy (anonymous):

Where are the asymptotes of f(x) = tan 4x from x = 0 to x = pi over two? @aum can you check this one please

OpenStudy (freckles):

Try writing tan in terms of sin and cos

OpenStudy (freckles):

That is a first step

OpenStudy (freckles):

once you have it in fractional form the equation you need to solve is bottom=0 for your asymptotes

OpenStudy (anonymous):

No the final answer @freckles

OpenStudy (freckles):

I'm not giving you the answer. You must get the answer. I'm showing you what to do (or telling you what to do). Have you wrote it in terms of sin and cos yet?

OpenStudy (anonymous):

I know the answer I just want to know what I did was correct

OpenStudy (freckles):

ok then show me

OpenStudy (anonymous):

I follow the same steps you given above and got x = pi over 4, x = 3 pi over 4 @freckles

OpenStudy (anonymous):

@freckles u coming

OpenStudy (freckles):

you are looking for when cos(u) is 0? because cos(pi) is -1 not 0

OpenStudy (aum):

tan 4x = sin 4x / cos 4x Vertical asymptotes exist when the denominator cos(4x) is zero. In the interval [0, 2pi], cos(z) = 0 when z = pi/2 or z = 3pi/2 Therefore, cos(4x) = 0 when 4x = pi/2 or 4x = 3pi/2 x = pi/8 or x = 3pi/8. Both these x values are within the specified interval of [0. pi/2]